OpenStudy (anonymous):

–2 * [–10.5 ÷ (2 + 3)]

OpenStudy (lalipapa123):

do you have any idea?

OpenStudy (igreen):

PEMDAS, first simplify the parenthesis. 2 + 3 = ?

OpenStudy (anonymous):

no I have two calulators telling me all diffrent things

OpenStudy (anonymous):

that's 5 iGreen

OpenStudy (lalipapa123):

first add 2+3 then divide it with -10.5 then muti ply the answer with -2

OpenStudy (igreen):

Yes, so we have: \(\sf -2 \times [-10.5 \div 5]\) Now divide -10.5 / 5

OpenStudy (anonymous):

-2.1

OpenStudy (lalipapa123):

u will get -4.2

OpenStudy (anonymous):

one says that and one says 4.2 plain

OpenStudy (igreen):

Yes, now multiply by -2.

OpenStudy (igreen):

-2 * -2.1 = ?

OpenStudy (lalipapa123):

sory its 4.2 because there is (-) sign

OpenStudy (lalipapa123):

(-) *(-) = +

OpenStudy (lalipapa123):

if my answer helped you plzzz medal me

OpenStudy (anonymous):

4.2

OpenStudy (lalipapa123):

yes!

OpenStudy (anonymous):

you give iGreen one then I will give you one medal

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (lalipapa123):

kk @coolman500

OpenStudy (lalipapa123):

i gave

OpenStudy (anonymous):

and I gave

OpenStudy (lalipapa123):

thanks. any other question?

OpenStudy (anonymous):

ya one sec

OpenStudy (lalipapa123):

kk..............